BMX Company has one employee. FICA Social Security taxes are 6.2% of the first $118,500 paid to its employee, and FICA MedicaRequired information (The following information applies to the questions displayed below.) BMX Company has one employee. FICAView transaction list Journal entry worksheet Prepare the employers September 30 journal entry to record the accrued salaryRequired information (The following information applies to the questions displayed below.) BMX Company has one employee. FICAView transaction list Journal entry worksheet Record the employers September 30 payroll taxes expense and its related liabil
Answer #1
a)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 2400 6.20% 148.80
FICA-Medicare 2400 1.45% 34.80
FUTA 1800 0.60% 10.80
SUTA 1800 2.90% 52.20
b)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 2800 6.20% 173.60
FICA-Medicare 2800 1.45% 40.60
FUTA 0 0.60% 0.00
SUTA 0 2.90% 0.00
c)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 5600 6.20% 347.20
FICA-Medicare 8700 1.45% 126.15
FUTA 0 0.60% 0.00
SUTA 0 2.90% 0.00
Note: Earnings taxable for FUTA and SUTA in (a) = 7000-5200 = $1800
Note: Earnings taxable for Social Security in (c) = 118500-112900 = $5600
2
Taxes to be Withheld From Gross Pay (Employee-Paid Taxes)
September Earnings Subject to Tax Tax Rate Tax Amount
Federal income tax 90.00
FICA—Social Security 2400 6.20% 148.80
FICA—Medicare 2400 1.45% 34.80
Total taxes withheld 273.60
Date General Journal Debit Credit
Sep 30 Salaries expense 2400.00
FICA—Social Security taxes payable 148.80
FICA—Medicare taxes payable 34.80
Employee federal income taxes payable 90.00
Salaries payable 2126.40
3
Employer Payroll taxes September earnings subject to tax Tax Rate Tax Amount
FICA-Social Security 2400 6.20% 148.80
FICA-Medicare 2400 1.45% 34.80
FUTA 1800 0.60% 10.80
SUTA 1800 2.90% 52.20
Total Payroll Taxes Expense 246.60
Date General Journal Debit Credit
Sep. 30 Payroll taxes expense 246.60
FICA—Social security taxes payable 148.80
FICA—Medicare taxes payable 34.80
Federal unemployment taxes payable 10.80
State unemployment taxes payable 52.20